Here is Venn diagram (figure 1 from ABVenn.mp): This modification (figure 2) adds color and outlines the figure better: And this modification (figure 3) makes simulated cross-hatching for the AB intersection. It also uses LaTeX commands to center and enlarged it.

This last figure shows how the cross-hatched intersection is created by skipping the “clip” command:

2